express each percent as a fraction or a mixed number in the simplest form.
a.)125% b.)580%
I know how to do this if the numbers are below 100, but I'm not for sure how to do this if they are over 100
Thank you

Found 2 solutions by stanbon, JLB:
Answer by stanbon(75887)

express each percent as a fraction or a mixed number in the simplest form.
---
% means "over 100".
a.)125% = 125/100 = 1.25
==========================
b.)580% = 580/100 = 5.8
============================
Cheers,
Stan H.

Answer by JLB(8)

1. For a percentage over 100%, take the first digit (1 in 125%) (5 in 580%) and make it the whole number.
1 ____________________ 5
2. Take the last two digits and make them the fractions (25/100 in 125%) (80/100 in 580%).
1 25/100 _____________ 5 80/100
3. Reduce the fractions (25/100 is 1/4) (80/100 is 4/5)
1 1/4 ____________ 5 4/5